OpenStudy (anonymous):

can someone help. If a pool deck of 650 square feet is to be laid with concrete 3" thick, what volume of concrete will have to be poured in cubic yards? (Give answer in nearest whole cubic yard.) ___ cubic yards

OpenStudy (anonymous):

What is the thickness of concrete? I didn't get that symbol "

OpenStudy (anonymous):

3"

OpenStudy (anonymous):

inches????

OpenStudy (anonymous):

I think its feet..inches correspond to symbol '

OpenStudy (anonymous):

the volume of concrete to be poured = Area of deck x thickness of concrete

OpenStudy (anonymous):

659x3/12

OpenStudy (anonymous):

650

OpenStudy (anonymous):

If the thickness is 3feet, the volume = 650 x 3 =1950 cubic feet = 650 cubic yards

OpenStudy (anonymous):

so that's not yet

OpenStudy (anonymous):

What else is in that?

OpenStudy (anonymous):

12 in =1 foot so we 3/12=0.5

OpenStudy (anonymous):

yes but 3/12=0.25

OpenStudy (anonymous):

1foot=12inches 1yard=3feet This is what that matters in this question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you want any help in solving it? And as far as i know, 3" = 3feet

OpenStudy (anonymous):

yes i need help im not sure

OpenStudy (anonymous):

so we have 650 right

OpenStudy (anonymous):

yes, the solution is exactly what i gave you earlier in a comment.

OpenStudy (anonymous):

Any doubt? Ask without any hesitation

OpenStudy (anonymous):

650/3???

OpenStudy (anonymous):

I will repeat the solution clearly..

OpenStudy (anonymous):

12 in=1 foot 1 yard=3 feet right

OpenStudy (anonymous):

Volume of concrete(in cubic feet) = Area of deck(in square feet) x thickness of concrete(in feet) = 650 x 3 = 1950 cubic feet 1yard = 3feet....so, 1foot = 1/3yard 1cubic feet = 1/27cubic yards 1950cubic feet = 1950/27 cubic yard = 72.22 cubic yards Sorry for the wrong answer i gave a few minutes ago.

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

they got 6 cubic yards i don't know how they got that

OpenStudy (anonymous):

okay then you are right about the units of thickness..it is inches.. Volume = 650 x 3/12 = 1950/12 cubic feet 1950/12 cubic feet = 1950/(12)(27) cubic yards = 72.22/12 cubic yards = 6 cubic yards approx.

OpenStudy (anonymous):

I kept 12 in the denominator throughout the process just to make it simple for you to understand the difference between the previous solution and the present one.